OpenStudy (anonymous):

One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side. One side of a square is 5n inches long. Another side is n + 16 inches long. Write an equation by setting these equal and solving for n. How long is a side of the square? One point for finding n and one point for finding the length of the side. @Mathematics

jimthompson5910 (jim_thompson5910):

Hint: all sides of a square are equal So if one side is n+16 inches long and another is 5n inches long, then n+16 = 5n Solve this equation to find n. Then use this value of n to find the length of the side.

OpenStudy (anonymous):

So will the answer be 1?

jimthompson5910 (jim_thompson5910):

you mean n = 1?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

no, n+16 = 5n 16 = 5n-n 16 = 4n 16/4 = n 4 = n n = 4

jimthompson5910 (jim_thompson5910):

so the side length is n+16 = 4+16 = 20 inches Notice that 5n = 5*4 = 20 also, so this confirms the answer

OpenStudy (anonymous):

So basically all you have to do to find the answer to one of the sides is just substitute and add?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

okay thanks for all your help i think i get how to do it now. :)